# Splat-Nav: Safe Real-Time Robot Navigation in Gaussian Splatting Maps

### v1

@@ Abstract
We present Splat-Nav, a navigation pipeline that consists of a real-time safe planning module and a robust state estimation module designed to operate in the Gaussian Splatting (GSplat) environment representation, a popular emerging 3D scene representation from computer vision. We formulate rigorous collision constraints that can be computed quickly to build a guaranteed-safe polytope corridor through the map. We then optimize a B-spline trajectory through this corridor. We also develop a real-time, robust state estimation module by interpreting the GSplat representation as a point cloud. The module enables the robot to localize its global pose with zero prior knowledge from RGB-D images using point cloud alignment, and then track its own pose as it moves through the scene from RGB images using image-to-point cloud localization. All of these modules operate mainly on CPU, freeing up GPU resources for tasks like real-time scene reconstruction. We demonstrate the safety and robustness of our pipeline in both simulation and hardware, where we show re-planning at 5 Hz and pose estimation at 20 Hz, an order of magnitude faster than Neural Radiance Field (NeRF)-based navigation methods, thereby enabling real-time navigation.
@@ I Introduction
Constructing a representation of the environment from data collected by onboard sensors, such as RGB(-D) cameras, lidar, or touch sensors, is a fundamental challenge for any autonomous system. Traditionally, these representations have included (watertight) triangular meshes [7], occupancy grids [8], point clouds [21], and Signed Distance Fields (SDFs) [33]. In each of these situations, the geometry is well-defined and there are many methods for planning and localizing within these maps.
Neural Radiance Fields (NeRFs) [30] have recently emerged in the computer vision community as a promising 3D scene representation. NeRFs offer several potential benefits over traditional scene representations: they provide photo-realistic renderings, they can be trained using only monocular RGB images, and current implementations can train NeRFs in seconds [31, 43]. However, NeRFs implicitly represent the environment as a volumetric density using a multilayer perceptron (MLP) deep neural network. This makes scene rendering of the environment computationally expensive and makes it difficult to use NeRFs for standard robotics tasks in an interpretable way, especially for tasks that require interacting with the environment.
More recently, Gaussian Splatting (GSplat) [20] has emerged as an expressive scene representation and differentiable rendering pipeline, trained from RGB images. GSplat represents the environment as a union of ellipsoidal objects, learning the number of ellipsoids along with the position, extents, color, and opacity of each. This approach allows for faster reconstruction by taking advantage of fast differential rasterization techniques for 3D Gaussians on the GPU and often yields more photorealistic renderings compared to NeRFs. More importantly for robotics, the use of (convex) geometric primitives facilitates the development of rigorous motion planning algorithms that are safe, robust, and real-time.
@@
Figure 1: Our real-time, robust navigation pipeline, Splat-Nav, consists of a safe planning module, Splat-Plan, and robust localization module, Splat-Loc, both powered by the Gaussian Splatting environment representation. We first represent our agent as an ellipsoid, and then rapidly solve for smooth paths using an ellipsoid-to-ellipsoid collision test. We then use the point cloud representation from the Gaussian Splat map to perform both global camera localization from RGB(-D) images using point cloud alignment, and online pose tracking using only RGB images with a PnP image-to-point cloud pose refinement. Additionally, we also learn a semantic channel to direct the robot towards regions of space that benefit the localization.
In this paper, we develop a framework called Splat-Nav, outlined in Fig. 1, for robot trajectory planning and pose estimation that is intertwined with the GSplat representation. The first component, Splat-Plan, generates safe trajectories through the GSplat scene. The primary contribution is a computationally efficient collision checker that is guaranteed to be safe. We demonstrate the flexibility of this checker by using it within two different motion planners: a sampling-based planner (RRT) [24] and a spline generation method [27]. We believe the theory presented in this work may also be useful for other planning methods as well as downstream applications built on Gaussian Splatting.
The second component, Splat-Loc, allows the robot to localize itself within the GSplat map using images from its onboard camera. The primary contribution is interpreting the GSplat map as a point cloud with features including the color and opacity values. Using this observation, we can use robust, off-the-shelf point cloud alignment algorithms to perform global pose alignment with zero prior pose knowledge (i.e., the kidnapped robot problem). We then perform real-time local image-to-point cloud pose estimation for the robot to track its motion through the scene, leveraging the high rendering speed of Gaussian Splatting. Both these modules are many times faster than their existing neural network-based NeRF counterparts, which do not provide guarantees on performance [1].
@@
To showcase these two modules, we provide simulation and hardware results to show that our pipeline: (1) generates safe, but not conservative, trajectories through the environment, (2) robustly localizes a robot in the scene, and (3) can be executed in real-time. We contrast our paths to those generated from prior work [1] that used NeRFs as the scene representation and to other baselines. Our planner demonstrates generalizable safe behavior and is not conservative. We also contrast localization results from the same work. We find that our localization is more accurate, much faster, and fails less often. We demonstrate re-planning at 5 Hz and online pose estimation at 20 Hz on a desktop computer, enabling real-time navigation.
@@ II-A1 Planning
There is a long history of planning within robotics. LaValle [22] provides an excellent explanation of the major algorithms in this realm. Most relevant for this work are the graph-based planners (e.g., A), which compute a path over a grid representation of the environment; sampling-based planners (e.g., PRM [19], RRT [23], and RRT [17]), which generate a path by sampling candidate states within the configuration space of the robot; and trajectory optimization-based planners (e.g. CHOMP [35] and Traj-Opt [39]), which take an optimization-based approach to planning. The Open Motion Planning Library (OMPL) [40] provides open-source implementations of many of these algorithms, which we use later for both path initialization and benchmarking. Liu et al. [27] utilizes an optimization-based approach in path planning, taking a point cloud representation of the environment and converting this into a set of safe polytopes. The resulting safe polytopes are utilized in computing a safe trajectory, parameterized as a spline, from a quadratic program (QP), which can be efficiently solved.
There is also extensive literature on planning based on onboard sensing. Typically, these works present reactive control schemes [48, 49, 50], using the sensed depth directly to perform collision checking in real-time. These methods typically are myopic, reasoning only locally about the scene. Consequently, such methods often converge to local optima in many problems, hindering the robot from reaching its goal, especially in cluttered/intricate environments. An alternate approach is to construct a map of the environment using the depth measurements. Often, a Signed Distance Field (SDF) or its truncated variant (TSDF) is constructed from depth data [32, 13], which is often encoded within a voxel representation. Such a representation is typical in dynamic robotic motion planning, providing fast collision checking and gradients in planning.
@@ II-A2 Localization
Prior work in robot localization utilize filtering schemes, such as Extended Kalman Filters (EKFs) [12, 9], Particle Filters (PFs) [10, 54], and other related filters [46, 2], to solve the pose localization problem. These methods generally estimate the pose of the robot from low-dimensional observations (measurements), extracted from the high-dimensional observations collected by the robot’s onboard sensors, such as cameras. This approach often fails to leverage the entire information available in the raw, high-dimensional measurements. Learning-based filtering methods [18, 16, 42] seek to address this limitation using deep-learning to develop end-to-end frameworks for localization, computing a pose estimate directly from raw camera images. Although learning-based approaches can be quite effective given sufficient training data, these methods are often limited to a single robot platform (dynamics model), and thus require separate filters for each robot or environment.
@@ II-B1 Planning
NeRFs implicitly represent the environment as a spatial density field (with color) [30]. There are a number of works on planning in this representation. Adamkiewicz et al. [1] plan trajectories for differentially flat robots, such as quadrotors, that minimize a total collision cost. Chen et al. [3] convert the NeRF into a probabilistic voxel grid and then use this to generate trajectories parameterized as Bézier curves. Tong et al. [45] uses the predicted depth map at sampled poses to enforce step-wise safety using a control barrier function. The above works are complementary, with Adamkiewicz et al. [1], Chen et al. [3] serving as high-level planners that encourage non-myopic behavior while Tong et al. [45] can be used as a safety filter for a myopic low-level controller.
@@ II-B2 Localization
There is some existing work on tracking the pose of a robot equipped with an on-board camera and IMU through a pre-trained NeRF map. iNeRF [52] does this for single images, and NeRF-Nav [1] and Loc-NeRF [28] both track a trajectory using a sequence of images. Other works consider simultaneous localization and mapping (SLAM) using a NeRF map representation. Sucar et al. [41], Yen-Chen et al. [52], Zhu et al. [55] all simultaneously optimize the NeRF weights and the robot/camera poses. Rosinol et al. [36] propose a combination of an existing visual odometry pipeline for camera trajectory estimation together with online NeRF training for the 3D scene.
@@ II-C GSplat Maps
To the best of our knowledge, there are no planning algorithms designed to operate in GSplat environments. There are a few recent works on SLAM using a GSplat representation of the environment [51, 53, 29]. However, none of these works have a standalone real-time pose localization component.
III Planning
@@
Now we present our planning pipeline based on Gaussian Splatting, called Splat-Plan. The pipeline consists of generating safe polytopic corridors (inspired by [27]) that discretizes the free space of a GSplat, rigorously built on theory derived from tests for intersection between ellipsoids. The method is also fast enough to provide real-time operation, provides generalizable safety guarantees, and is not conservative.
III-A Problem Statement
Our approach leverages the ellipsoidal representation of objects in a GSplat. An ellipsoid $\mathcal{E}_{j}$ is parameterized by covariance $\Sigma_{j}$ and mean $\mu_{j}$:
$\mathcal{E}_{j}=\{x\in\mathbb{R}^{3}\mid(x-\mu_{j})^{T}\Sigma_{j}^{-1}(x-\mu_{j})\leq 1\}.$ (1)
Given a bounding ellipsoid $\mathcal{E}_{\mathcal{R}}$ for the robot and a union of $N$ ellipsoids $\mathcal{G}=\{\mathcal{E}_{j}\}^{N}_{j=1}$ (from GSplat), we seek to find a feasible path from an initial position $X$ to a goal position $Y$ such that there are no collisions, i.e., $\mathcal{E}_{\mathcal{R}}\cap\mathcal{E}_{j}=\emptyset,\,\forall\mathcal{E}_{j}\in\mathcal{G}$.
III-B Minimal Bounding Ellipsoid
Typically, robots are represented by a point cloud or a mesh. A mesh of $M$ vertices can be converted to a point cloud by selecting its vertices to yield $\mathcal{P}_{\mathcal{R}}=\{x_{i}\}_{i=1}^{M}$. In order to convert these canonical representations to the one we desire, we can solve a simple convex program to generate our ellipsoid $\mathcal{E}_{\mathcal{R}}$. To do this, we solve the following convex program:
@@ III-B Minimal Bounding Ellipsoid
Then,the mean of $\mathcal{E}_{\mathcal{R}}$ is given by $\mu_{\mathcal{R}}=-A^{-1}b$ and the covariance $\Sigma_{\mathcal{R}}=(A^{T}A)^{-1}$. Note, if the robot body consists of many points, computing the convex hull of these points first, and using the convex hull vertices in (2) will drastically lower the computation costs.
III-C Collision Detection
The environment and the robot body are represented completely by ellipsoids or unions thereof. Therefore, collision is identical to finding intersection between two ellipsoids (the robot body and an ellipsoid $\mathcal{E}_{j}$ from the environment). This is given by the following Theorem.

### v2

@@ Abstract
We present Splat-Nav, a real-time navigation pipeline designed to work with environment representations generated by Gaussian Splatting (GSplat), a popular emerging 3D scene representation from computer vision. Splat-Nav consists of two components: 1) Splat-Plan, a safe planning module, and 2) Splat-Loc, a robust pose estimation module. Splat-Plan builds a safe-by-construction polytope corridor through the map based on mathematically rigorous collision constraints and then constructs a Bézier curve trajectory through this corridor. Splat-Loc provides a robust state estimation module, leveraging the point-cloud representation inherent in GSplat scenes for global pose initialization, in the absence of prior knowledge, and recursive real-time pose localization, given only RGB images. The most compute-intensive procedures in our navigation pipeline, such as the computation of the Bézier trajectories and the pose optimization problem run primarily on the CPU, freeing up GPU resources for GPU-intensive tasks, such as online training of Gaussian Splats. We demonstrate the safety and robustness of our pipeline in both simulation and hardware experiments, where we show online re-planning at 5 Hz and pose estimation at about 25 Hz, an order of magnitude faster than Neural Radiance Field (NeRF)-based navigation methods, thereby enabling real-time navigation. Videos are hosted on our website, and code can be found here.
@@ I Introduction
Autonomous robotic operation requires robots to plan safe paths to reach a desired goal location as well as to localize themselves during navigation to perform closed-loop control to follow the planned path. These two areas of planning and localization are core to mobile robotics and robotic manipulation, and the specific solutions for each problem depend upon the representation used for the map of the environment. Traditionally, these representations have included occupancy grids [1], (watertight) triangular meshes [2], point clouds [3], and Signed Distance Fields (SDFs) [4], all of which provide well-defined geometry. Many planning and localization algorithms have been designed for these representations (see Section II-A). However, these explicit scene representations are generally constructed at limited resolutions, to enable real-time operation, leaving out potentially-important scene details that could be valuable in planning and localization problems.
@@
Figure 1: Our real-time, robust navigation pipeline, Splat-Nav, consists of a safe planning module, Splat-Plan, and robust localization module, Splat-Loc, both powered by the Gaussian Splatting environment representation. We first represent our agent as an ellipsoid, and then rapidly solve for smooth paths using an ellipsoid-to-ellipsoid collision test. We then use the point cloud representation from the Gaussian Splat map to perform both global camera localization from RGB(-D) images using point cloud alignment, and online pose tracking using only RGB images with a PnP image-to-point cloud pose refinement. Additionally, we also learn a semantic channel to enable semantically-conditioned goal locations like “go to the table“.
Implicit scene representations such as Neural Radiance Fields (NeRFs) [5] address these fundamental limitations. NeRFs, which were introduced by the computer vision community, represents the environment as a volumetric density and view-dependent color using a multilayer perceptron (MLP) deep neural network. NeRFs offer several potential benefits over traditional scene representations: they provide more photo-realistic renderings, they can be trained using only monocular RGB images, and current implementations of NeRFs can be trained in seconds [6, 7]. Although some NeRF methods can be trained quickly, these methods generally require access to high-performance GPUs, which might not be available in many robotics tasks. We note that some planning and localization algorithms for NeRF maps exist (see Section II-B); however, these methods do not run in real-time, in general. Moreover, the use of MLPs makes scene rendering computationally expensive and makes it difficult to use NeRFs for standard robotics tasks in an interpretable way, especially for tasks that require interacting with the environment.
More recently, Gaussian Splatting (GSplat) [8] has emerged as a viable scene representation compared to NeRFs, representing the environment as a collection of Gaussians associated with a position (mean), spatial extents, color, and opacity. Compared to NeRFs, GSplat generates more photorealistic renderings, has shorter or comparable training times, and has much faster rendering by taking advantage of differentiable rasterization techniques for 3D Gaussians on the GPU. More importantly for robotics, GSplats, unlike NeRFs, offer a geometrically consistent collision geometry, enabling us to use level sets of these Gaussians to generate an ellipsoidal representation of the scene (see Section III for more details). These interpretable geometric primitives facilitate the development of rigorous motion planning algorithms that are safe, robust, and real-time.
I-A Approach
In this paper, we develop a framework called Splat-Nav, outlined in Fig. 1, for robot trajectory planning and pose estimation that is designed to work with the GSplat representation. The first component, Splat-Plan, generates safe trajectories through the GSplat scene. The primary contribution is an efficient navigation pipeline for Gaussian Splatting Maps, built upon a computationally efficient collision checker (which harnesses GPU parallelization to enable real-time collision checking) and a safe planning framework. The collision checker extends the method in [9] for ellipsoid-to-ellipsoid collision checking by eliminating computationally-intensive root-finding procedures required in the method through the formulation of a generalized eigenvalue problem to yield a significant computational speed-up. We implement the collision-checking algorithm with efficient GPU parallelization using sampling to obtain massive speedups for checking millions of ellipsoids collisions per second. Moreover, we integrate safe planning tools into our proposed planning pipeline through the computation of safe polytopes obtained from the decomposition of the configuration space into occupied and collision-free regions. We compute Bézier spline trajectories for the robot within these safe polytopes for navigation.
The second component, Splat-Loc, allows the robot to localize itself within the GSplat map using monocular images from its onboard camera. The primary contribution is the derivation of a real-time pose estimation module, which leverages the real-time rendering speeds of GSplats for fast extraction of point clouds for image-based localization. We introduce a robust image-to-point cloud pose estimation pipeline, built on well-established computer vision tools, harnessing the visual and geometric features of images to compute an estimate of the robot’s pose, without having to rely on more traditional gradient-based methods. In addition, we present our pose estimator within an optimization-based paradigm, enabling the seamless integration of prior knowledge from the planning module, as well as scene-relevant constraints. Moreover, in the absence of a prior guess on the pose of the robot, we provide a global pose estimation module, utilizing point-cloud registration algorithms, provided an RGB-D camera is available. Both modules are many times faster than their existing NeRF-based counterparts, which do not provide guarantees on performance [10].
@@ I-A Approach
To showcase these two modules, we provide simulation and hardware results to show that our pipeline: (1) generates safe, but not overly conservative, trajectories through the environment, (2) robustly localizes a robot in the scene, and (3) can be executed in real-time. Our planner demonstrates safe flight over a variety of synthetic and real-world GSplat maps. We find that our localization is more accurate, faster, and fails less often compared to baselines. We demonstrate re-planning at 5 Hz and online pose estimation at about 25 Hz on a desktop computer, enabling real-time navigation.
I-B Contributions
The key contributions of this paper are as follows: 1. We derive an accelerated collision checking algorithm for the potentially millions of ellipsoids that make up GSplat scenes, parallelized on a GPU for real-time collision checking (Proposition 1, Corollary 1, Corollary 2). 2. We generate safe polytope corridors from the ellipsoidal representation of GSplat maps (Proposition 2). 3. We implement a receding horizon trajectory planning algorithm based on the flight corridors (Sec. IV-D). 4. We propose a global pose initialization and a UKF-based filter for state estimation from on-board vision, by adapting existing point cloud alignment and Perspective-$n$-Point (PnP) algorithms to the GSplat representation (Sec. V). 5. We show significantly improved computational speed and lower tracking errors in comparison to existing state estimation algorithms for NeRFs and GSplat maps (Tables I and II, and Fig. 7). 6. We demonstrate real-time operation in hardware tests on a quadrotor drone (Sec. VI).
@@ II-A1 Planning
There is a long history of planning within robotics. We refer readers to [11] for an excellent explanation of the major algorithms in this realm. Most relevant for this work are the graph-based planners (e.g., A), which compute a path over a grid representation of the environment; sampling-based planners (e.g., PRM [12], RRT and RRT [13]), which generate a path by sampling candidate states within the configuration space of the robot; and trajectory optimization-based planners (e.g. CHOMP [14] and Traj-Opt [15]), which take an optimization-based approach to planning. The Open Motion Planning Library (OMPL) [16] provides open-source implementations of many of these algorithms, which we use later for benchmarking. Prior work in [17] utilizes an optimization-based approach in path planning, taking a point cloud representation of the environment and converting this into a set of safe polytopes. The resulting safe polytopes are utilized in computing a safe trajectory, parameterized as a spline from a quadratic program (QP), which can be efficiently solved.
There is also extensive literature on planning based on onboard sensing. Typically, these works present reactive control schemes [18, 19], using the sensed depth directly to perform collision checking in real-time. These methods typically are myopic, reasoning only locally about the scene. Consequently, such methods often converge to local optima, preventing the robot from reaching its goal, especially in cluttered/intricate environments. An alternate approach is to construct a map of the environment using the depth measurements. Often, a Signed Distance Field (SDF) or its truncated variant (TSDF) is constructed from depth data [20, 21], which is often encoded within a voxel representation. Such a representation is typical in dynamic robotic motion planning, providing fast collision checking and gradients in planning.
@@ II-A2 Localization
Prior works in robot localization utilize filtering schemes, such as Extended Kalman Filters (EKFs) [22, 23], Particle Filters (PFs) [24, 25], and other related filters [26, 27], to solve the pose localization problem. These methods generally estimate the pose of the robot from low-dimensional observations (measurements), extracted from the high-dimensional observations collected by the robot’s onboard sensors, such as cameras. This approach often fails to leverage the entire information available in the raw, high-dimensional measurements. Learning-based filtering methods [28, 29] seek to address this limitation using deep-learning to develop end-to-end frameworks for localization, computing a pose estimate directly from raw camera images. Although learning-based approaches can be quite effective given sufficient training data, these methods are often limited to a single robot platform (dynamics model), and thus require separate filters for each robot or environment.
@@ II-B1 Planning
NeRFs implicitly represent the environment as a spatial density field (with color) [5]. Using this representation, NeRF-Nav [10] plans trajectories for differentially flat robots, e.g., quadrotors, that minimize a total collision cost. Further, CATNIPS [30] converts the NeRF into a probabilistic voxel grid and then uses this to generate trajectories parameterized as Bézier curves. The work in [31] uses the predicted depth map at sampled poses to enforce step-wise safety using a control barrier function. The above works are complementary, with [10, 30] serving as high-level planners that encourage non-myopic behavior while [31] can be used as a safety filter for a myopic low-level controller.
@@ II-B2 Localization
There is some existing work on tracking the pose of a robot equipped with an on-board camera and IMU through a pre-trained NeRF map. iNeRF [32] does this for single images, and NeRF-Nav [10] and Loc-NeRF [33] both track a trajectory using a sequence of images. Other works consider simultaneous localization and mapping (SLAM) using a NeRF map representation. Existing methods such as [32, 34] all simultaneously optimize the NeRF weights and the robot/camera poses. NeRF-SLAM [35] proposes a combination of an existing visual odometry pipeline for camera trajectory estimation together with online NeRF training for the 3D scene.
@@ II-C GSplat Maps
To the best of our knowledge, there are no planning algorithms designed to operate in GSplat environments. There are a few recent works on SLAM using a GSplat representation of the environment [36, 37, 38]. However, none of these works have a standalone real-time pose localization component, and none address safe trajectory planning or control through the GSplat map, as is our focus in this paper.
III 3D Gaussian Splatting
We present a brief introduction to 33D Gaussian Splatting [8], a radiance field method for deriving volumetric scene representations. In contrast to explicit 3D scene representations (e.g., point clouds and meshes) and continuous scene representations (e.g., NeRFs), Gaussian Splatting represents non-empty space in a scene using 33D Gaussian primitives, each of which is parameterized by a mean $\mu\in\mathbb{R}^{3}$} (defining its position), covariance matrix $\Sigma\in\mathbb{S}_{++}$} (related to its spatial extent), opacity $\alpha\in[0,1]$}, and spherical harmonics (SH) coefficients (defining view-dependent colors). For better numerical optimization, the anisotropic 33D covariance of each Gaussian is written as: $\Sigma=RSS^{T}R^{T}$}, where $R\in\mathrm{SO}(3)$} is a rotation matrix (parameterized by a quaternion) and $S$ is a diagonal scaling matrix (parameterized by a 33D vector). The number of primitives, along with the coefficients for each primitive, are learned using monocular images (the same as NeRF methods) along with a sparse point-cloud representation of the environment, which is generally computed from structure-from-motion [39].
The anisotropic covariance of the Gaussian primitives along with adaptive density control enable the computation of compact high-quality representations, even in complex scenes, unlike many state-of-the-art point-based rendering methods. Further, 33D Gaussian Splatting obviates the need for volumetric ray-marching required in NeRF methods, enabling high-quality real-time rendering, even from novel views. Moreover, 33D Gaussian Splatting enables relatively fast extraction of a mesh representation of the scene from the Gaussian primitives. In Fig. 2, we visualize the ground-truth mesh, the GSplat mesh, and the associated point cloud extracted from a NeRF of a simulated Stonehenge scene. We note that the resulting GSplat mesh has a smaller Chamfer distance (0.031 with 3M vertices) compared to the NeRF point cloud (0.081 with 4M points) despite having fewer points. We note that the NeRF does not necessarily yield a view-consistent geometry due to volumetric rendering, especially when the point cloud is not post-processed to remove outliers.
To render an image from a given camera pose, the $3D$ Gaussians are projected onto the image plane using an affine approximation of the projective transformation, given by $\Sigma_{2D}=JW\Sigma W^{T}J^{T}$}, with Jacobian $J$ and viewing transformation $W$. Subsequently, the color $C$ of each pixel is computed via $\alpha$-blending of the splats, given by:
$C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
where $c_{i}$ represents the color of Gaussian $i$, computed from the Spherical Harmonics model, $\mathcal{N}$} represents the set of Gaussians whose projection intersects with the pixel. We overload the definition of $\alpha_{i}$ in the rendering model to represent the opacity of Gaussian $i$ multiplied by the unnormalized probability density function associated with each Gaussian in the 22D image plane, given by: $G(x)=e^{-\frac{1}{2}(x-\mu)^{T}\Sigma_{2D}^{-1}(x-\mu)},$} with $\Sigma_{2D}\in\mathbb{S}_{++}$} representing the $2D$ covariance matrix. We note the rendering model (1) corresponds to the rendering model used in NeRF methods [5], although Gaussian Splatting utilizes differentiable tile-based rasterization, contributing to its high rendering speed.
Remark 1.
The original work only projects $3D$ Gaussians whose $99\%$ confidence interval intersects the view frustum of a camera, effectively restricting the scene representation to the $99\%$ confidence ellipsoid associated with each Gaussian. Consequently, the union of the $99\%$ confidence ellipsoids represents the entirety of the geometry of the scene learned during the training procedure.
The Gaussian Splatting representation is trained via stochastic gradient descent with a loss function comprising of the photometric loss between the rendered and ground-truth images and the structural-similarity (SSIM) index loss, while simultaneously adaptively controlling the number of Gaussians in the representation. Generally, Gaussian Splatting requires about the same training time as state-of-the-art NeRF methods, while achieving about the same or better quality. However, Gaussian Splatting achieves real-time rendering speeds [8], unlike NeRFs. See Fig. 2 for an example showing a NeRF and GSplat map of the same scene. In the subsequent sections, we present the core contributions of our work in deriving an efficient navigation pipeline for robots, describing how we leverage 33D Gaussian Splatting as the underlying scene representation, which we augment with semantic embeddings for goal specification in natural-language. Specifically, we present our planning and pose estimation pipelines that enable safe navigation in Gaussian Splatting environments.
Figure 2: Visualization of a point cloud from a NeRF and a mesh from a Gaussian Splat in the synthetic scene Stonehenge. The Chamfer Distance between the NeRF and ground-truth is 0.081 (with 4M points). The Chamfer distance between the GSplat and ground-truth is 0.031 (with 3M vertices). The collision geometry (especially the foreground) of the GSplat is better and can be extracted instantaneously from the model parameters compared to the costly rendering procedure from many viewpoints to create a point cloud from the NeRF.
IV Planning in Gaussian Splatting Maps
@@
Now, we present our planning pipeline based on Gaussian Splatting, called Splat-Plan. The pipeline consists of generating safe polytopic corridors (inspired by [17]) that discretize the free space of a GSplat Map, for navigation from an initial configuration to a goal configuration, rigorously built on theory derived from tests for intersection between ellipsoids. The method is fast enough to provide real-time operation, provides safety guarantees extending to any scene with a pre-trained GSplat representation, and is not overly-conservative. We stress that, as with any safety guarantee on a map, our ultimate safety rests on the completeness of the map. If the map does not reflect the presence of an obstacle, our method may collide with the obstacle—we cannot avoid what we cannot see. In practice, we observe that GSplat maps provide fast and efficient representations of the underlying ground truth geometry, as validated in our hardware experiments.
Before presenting the planning problem, we make the following assumptions on the representations of the robot $\mathcal{R}$ and the map $\mathcal{G}$ considered in this work. We assume that the robot is represented by a union of the ellipsoids in the non-empty set $\{\mathcal{E}_{\mathcal{R},i}\}_{i=1}^{d}$, where $d$ denotes the cardinality of the set, i.e., $\mathcal{R}\subseteq\cup_{i=1}^{d}\mathcal{E}_{\mathcal{R},i}$}. For simplicity, we consider a singleton set $\mathcal{E}_{\mathcal{R}}$, noting that the subsequent discussion applies directly to the non-singleton case by running the collision check for all robot ellipsoids. One can also convert a mesh or point cloud of a robot to an ellipsoid by finding the minimal bounding ellipsoid (or sphere).
We represent non-empty space in the environment with $\gamma\%$ confidence ellipsoids obtained from the GSplat map, as discussed in Remark 1, given by:
$\mathcal{E}_{j}=\{x\in\mathbb{R}^{3}\mid(x-\mu_{j})^{T}\Sigma_{j}^{-1}(x-\mu_{j})\leq\chi_{3}^{2}(\gamma)\},$ (2)
@@
Remark 2 (Online Gaussian Splatting).
Our planning algorithm requires a GSplat map. This map can be constructed a priori or online using real-time SLAM methods for radiance fields [38, 37, 36], which is a very new and active area of research.
Remark 3 (Handling Uncertainty of the Scene Representation).
We can vary the value of $\gamma$} (from that used during the training procedure) based on the quality of the GSplat Map and uncertainty in different regions of the Gsplat Map. In general, larger values of $\gamma$ inflate the volume of the confidence ellipsoids associated with each Gaussian, resulting in greater safety margins and more conservative planning. The converse holds if smaller values of $\gamma$ are selected. Moreover, for simplicity, we utilized a uniform value of $\gamma$. However, the value of $\gamma$ can vary among the ellipsoids, allowing the planner to account for varying levels of uncertainty in different regions of the GSplat Map. Likewise, the volume of the ellipsoid representing the robot can be increased/decreased to account for uncertainty in the pose of the robot.
Remark 4 (Dynamic Scenes).
We limit our discussion to planning in static scenes. However, we note that our method readily applies to planning in dynamic scenes, under the assumption that a dynamic Gaussian Splatting scene representation can be constructed. We discuss more about planning in dynamic scenes in Section VIII.
IV-A Problem Statement
Given a bounding ellipsoid $\mathcal{E}_{\mathcal{R}}$ for the robot and a GSplat Map $\mathcal{G}$, we seek to find a feasible path for a robot to navigate from an initial configuration $X$ to a specified goal configuration $Y$, such that there are no collisions, i.e., $\mathcal{E}_{\mathcal{R}}\cap\mathcal{E}_{j}=\emptyset,\,\forall\mathcal{E}_{j}\in\mathcal{G}$}.
IV-B Collision Detection
We leverage the ellipsoidal representations of the robot and the environment to derive an efficient collision-checking algorithm, based on [9], where we take advantage of GPU parallelization for faster computation. We build upon [9] rather than on other existing ellipsoid-to-ellipsoid intersection tests, because of its amenability to significant GPU parallelization. We do not utilize the GJK algorithm [40], since we do not require knowledge of the distance between the two ellipsoids. For completeness, we restate the collision-checking method from [9, Proposition 2].